18.- If three people, randomly selected, are arrested on the street, what are the chances that
a) all of them were born on Friday; A= 0.0029
b) two of them were born on Friday; A=0.0525
c) None of them was born on Monday? A= 0.6297
Answer #1

There are 7 days in a week.

So, P(Friday) = 1 / 7

a)

P(All three born on Friday) = 1/7 * 1/7 * 1/7

= 0.0029

b)

P(Friday) = 1/7 , P(Not Friday) = 1 - 1/7 = 6/7

P(Two of them born on Friday) = 3C2 * (1/7 * 1/7 *6/7)

= 3 * 0.01749

= 0.0525

c)

P(Monday) = 1 / 7 , P(Not Monday) = 1 - 1/7 = 6/7

P(None of them born on Monday) = 6/7 * 6/7 * 6/7

= 0.6297